OpenStudy (anonymous):

The perimeter of a rectangle is 42 centimeters. The length of the rectangle can be represented by x+4 and it's width can be represented by 2x-7. What are the dimensions of the rectangle in centimeters?

OpenStudy (anonymous):

Perimeter of rectangle is given by 2 ( l + w) 2(x + 4 + 2x - 7) = 42 2(3x - 3) = 42 2*3 (x - 1) = 42 x - 1 = 42/6 = 7 x = 8 Now Substitute that in length and width equations u get the answer

OpenStudy (anonymous):

Length x + 4 = 8 + 4 = 12cm and Width 2x - 7 = 2(8) - 7 = 16 - 7 = 9cm